# An Empirical Study of Validating Synthetic Data for Formula Generation

### v2

@@ Abstract
Large language models (LLMs) can be leveraged to help with writing formulas in spreadsheets, but resources on these formulas are scarce, impacting both the base performance of pre-trained models and limiting the ability to fine-tune them. Given a corpus of formulas, we can use a(nother) model to generate synthetic natural language utterances for fine-tuning. However, it is important to validate whether the NL generated by the LLM is indeed accurate to be beneficial for fine-tuning. In this paper, we provide empirical results on the impact of validating these synthetic training examples with surrogate objectives that evaluate the accuracy of the synthetic annotations. We demonstrate that validation improves performance over raw data across four models (2 open and 2 closed weight). Interestingly, we show that although validation tends to prune more challenging examples, it increases the complexity of problems that models can solve after being fine-tuned on validated data.
@@ 1 Introduction
Derived-column formulas in spreadsheets generate a new column by transforming existing columns in a table, and they have been shown challenging to write Gulwani et al. (2012). To aid users in writing such formulas, we can ask for a description in natural language Zhao et al. (2024). Unfortunately, since such formulas are sparse, therefore pre-trained language models (especially smaller) struggle in generating them without fine-tuning (0.03 for phi-2 in pass@10).
To construct a dataset for fine-tuning, public spreadsheet workbooks can be used but they contain only tables and formulas, whereas a fine-tuning dataset also requires paired natural language (NL) descriptions corresponding to each (Table, Formula). Traditionally datasets for NL-to-code tasks have been manually annotated Zhou et al. (2024); Austin et al. (2021). This is a time-consuming and expensive process. Leveraging LLMs, known for their text generation capabilities, is a viable alternative Tan et al. (2024) given that the synthetic NL generated by LLMs is accurate, as recent studies have shown that quality is more important than quantity Zhou et al. (2024); Li et al. (2023).
In this paper, we leverage LLMs to predict the accuracy of synthetic NL using 3 surrogate objectives, and show empirical results of fine-tuning models on subsets of synthetic data that are accepted by these objectives. Fine-tuning models on validated subsets shows better performance in predicting formulas compared to using raw data. For example, GPT-4 fine-tuned on data validated by alternative code generation objective saw up to a 25% improvement in evaluation scores along with a 23% reduction in training time. Additionally, we observe that the models fine-tuned on validated data perform better on more complex problems. Further, we release the synthetic dataset to seed future research in this area.
@@ 2 Related work
Tan et al. (2024) discusses the applications of LLMs in data annotation for classification tasks. Goel et al. (2023) demonstrates the use of LLMs in the medical domain, where they assist in labeling data with expert verification. Wang et al. (2024) and Kim et al. (2024) explore human-LLM collaborative approaches for annotation and verification. There has been no comparison of NL validation techniques on synthetic NL for NL2F.
@@ 3 Validating synthetic data
This validator asks the LLM to directly predict the output values $F(T)$ from $(\hat{U},T)$ and uses an element-wise row comparison to evaluate correctness. For numbers, we allow an absolute difference of 0.05. For strings, we use a longest common sub-sequence ratio of 0.8 as passing criterion. This approach leverages natural language to emulate the computation directly. It is inspired from the alternate task of output prediction discussed in Khatry et al. (2023)
@@ 4 Experimental setup
We use two open (Phi-2 and Mistral-7b-instruct) and two closed-weight (GPT-35-turbo and GPT-4) models. Phi-2 (8 $\times$ V100) and Mistral (1 $\times$ A100) were fine-tuned for 10 and 15 epochs respectively. We selected the best checkpoint using validation loss. GPT-35-turbo ($16$$\times$ A100) and GPT-4 ($24$$\times$ A100) were fine-tuned using the Azure API. Mistral, GPT-35-turbo, GPT-4 were fine-tuned using LoRA Hu et al. (2021).
@@
We use the pass@$k$ metric Chen et al. (2021a) based on execution match of formula, were $k$ represents the number of predictions considered out of the total number of predictions provided. In our evaluation system, we generate $n=10$ predictions at temperature $0.6$ and consider pass@1, pass@3 and pass@10.
@@ 5.1 RQ1: Comparing validators
Table 2: Performance comparison of the different models on SofSet Benchmark using pass@1, pass@3 and pass@10 metric. Three out of the four models give best performance when fine-tuned on data validated by $V_{P}$.
Phi-2 Mistral GPT-35-turbo GPT-4
FT on # ex P@1 P@3 P@10 P@1 P@3 P@10 P@1 P@3 P@10 P@1 P@3 P@10
Base 0 0.01 0.01 0.03 0.01 0.03 0.05 0.19 0.27 0.35 0.28 0.34 0.4
Raw 7833 0.05 0.07 0.10 0.09 0.14 0.20 0.22 0.28 0.32 0.25 0.30 0.36
$V_{O}$ 2266 0.03 0.05 0.07 0.11 0.15 0.19 0.21 0.27 0.29 0.27 0.32 0.37
$V_{P}$ 4095 0.06 0.08 0.12 0.08 0.12 0.16 0.23 0.28 0.33 0.31 0.38 0.45
$V_{C}$ 5246 0.06 0.08 0.10 0.09 0.13 0.16 0.23 0.28 0.33 0.31 0.37 0.44
Table 3: Pairwise comparison of performance of sub-sampled ($\subset$) data from validated ($V$) against rejected ($\neg V$) examples for different validation strategies on different models. The performance of $V_{O}$ is affected due to its poor handling of functions with numerical values.
Phi-2 Mistral GPT-35-turbo GPT-4
FT on # ex P@1 P@3 P@10 P@1 P@3 P@10 P@1 P@3 P@10 P@1 P@3 P@10
$V_{O}$ 2266 0.03 0.05 0.07 0.11 0.15 0.19 0.21 0.27 0.29 0.25 0.30 0.36
$\neg V_{O}$ 5567 0.03 0.04 0.06 0.10 0.14 0.18 0.21 0.26 0.29 0.26 0.33 0.37
$\subset{V_{C}}$ 2266 0.04 0.06 0.08 0.08 0.12 0.17 0.22 0.25 0.29 0.31 0.34 0.37
$\neg V_{C}$ 2587 0.02 0.03 0.05 0.07 0.10 0.15 0.21 0.24 0.28 0.28 0.32 0.36
$\subset V_{P}$ 2266 0.05 0.07 0.10 0.06 0.10 0.15 0.26 0.29 0.32 0.33 0.36 0.39
$\neg V_{P}$ 3738 0.03 0.04 0.07 0.06 0.10 0.13 0.18 0.23 0.27 0.27 0.31 0.34
Table 4: Training time for different models on data subsets. Models fine-tuned on $V_{P}$ and $V_{C}$ subsets require less time than on raw data while delivering better downstream performance.
Data Phi-2 Mistral GPT-35-turbo GPT-4
Raw 15h44m 8h51m 4h45m 14h00m
$V_{O}$ 4h17m 2h32m 1h54m 7h25m
$V_{P}$ 8h08m 4h50m 3h00m 10h50m
$V_{C}$ 10h3m 5h59m 3h50m 11h19m
@@ 5.2 RQ2: Effect on fine-tuning performance
Table 2 shows base model (few-shot) and fine-tuning performance on different subsets of data. For the smaller models, Phi-2 and Mistral, the performance increase with fine-tuning is more significant. Except for Mistral in pass@10 and GPT-35-turbo in pass@3, a smaller, validated dataset yields better performance than raw data. $V_{P}$ yields the best performance on average with nearly half the size of raw data. GPT-4 improves only when fine-tuned on validated data. Surprisingly, GPT-35-turbo without fine-tuning outperforms the fine-tuned version, likely due to differences in data distribution between training and testing benchmarks. Besides performance, fine-tuning with validated data also reduces training time significantly, as shown in Table 4.
@@
Table 3 compares the performance of fine-tuning on the accepted (& subsampled) and rejected ($\neg$) examples for each validator. We sub-sample the accepted sets to 2266—the number of examples in the smallest set ($V_{O}$). Results of pairs $(\subset V,\neg V)$ are marked in green if $(\subset V>\neg V)$, blue if $(\subset V=\neg V)$ and red if $(\subset V<\neg V)$. We observe that, despite the smaller size of the validated data subset (subsampled), it outperforms its larger invalidated (rejected) counterpart in most (28/36) comparisons. The only decrease in performance happens for $V_{O}$ on GPT-4, likely due to the many functions (51) that were eliminated from the training data.
@@ 5.3 RQ3: Analysing solved cases
Figure 3 shows properties of the solved cases (where pass@10 $=$ 1) after fine-tuning different models on raw data and validated subsets. We see that fine-tuning on datasets with fewer unique functions still enables all models (except for Mistral) to use more functions.
The average function call count increases for validated subsets compared to the raw data, indicating more complex formulas are solved by models fine-tuned on validated data. For GPT-4 and GPT-35-turbo, average operator count also increases with fine-tuning on validated data.
@@ 6 Conclusion
We empirically evaluate the effect of automated validation of synthetic data using LLMs on the fine-tuning performance of derived-column NL-to-formula. We validate synthetic NL annotations with three surrogate tasks (classification, code generation in Python, and output prediction) and fine-tune different models on the examples accepted by each of these methods. In general, fine-tuning on smaller, validated datasets improves performance. Despite validation resulting in datasets with simpler formulas, that does not cause the fine-tuned models to only solve simpler problems. Further, we release our dataset to seed research in this area.

### v3

@@ Abstract
Large language models (LLMs) can be leveraged to help with writing formulas in spreadsheets, but resources on these formulas are scarce, impacting both the base performance of pre-trained models and limiting the ability to fine-tune them. Given a corpus of formulas, we can use a(nother) model to generate synthetic natural language utterances for fine-tuning. However, it is important to validate whether the natural langauge (NL) generated by the LLM is indeed accurate to be beneficial for fine-tuning. In this paper, we provide empirical results on the impact of validating these synthetic training examples with surrogate objectives that evaluate the accuracy of the synthetic annotations. We demonstrate that validation improves performance over raw data across four models (2 open and 2 closed weight). Interestingly, we show that although validation tends to prune more challenging examples, it increases the complexity of problems that models can solve after being fine-tuned on validated data.
@@ 1 Introduction
Derived-column formulas in spreadsheets generate a new column by transforming existing columns in a table, and they have been shown challenging to write Gulwani et al. (2012). To aid users in writing such formulas, we can ask for a description in natural language Zhao et al. (2024). Unfortunately, since such formulas are sparse, therefore pre-trained language models (especially smaller) struggle in generating them without fine-tuning (for example, one of our models, Phi-2, achieved a pass@10 score of only 0.03, indicating a very low success rate in generating the correct formulas within 10 attempts.).
To construct a dataset for fine-tuning, public spreadsheet workbooks can be used but they contain only tables and formulas, whereas a fine-tuning dataset also requires paired natural language (NL) descriptions corresponding to each (Table, Formula). Traditionally datasets for NL-to-code tasks have been manually annotated Zhou et al. (2024); Austin et al. (2021). This is a time-consuming and expensive process. Leveraging LLMs, known for their text generation capabilities, is a viable alternative Tan et al. (2024) given that the synthetic NL generated by LLMs is accurate, as recent studies have shown that quality is more important than quantity Zhou et al. (2024); Li et al. (2023); lozhkov2024fineweb-edu.
In this paper, we leverage LLMs to predict the accuracy of synthetic NL using 3 surrogate objectives, and show empirical results of fine-tuning models on subsets of synthetic data that are accepted by these objectives. Fine-tuning models on validated subsets shows better performance in predicting formulas compared to using raw data. For example, GPT-4 fine-tuned on data validated by alternative code generation objective saw up to a 28% improvement in evaluation scores along with a 23% reduction in training time. Additionally, we observe that the models fine-tuned on validated data perform better on more complex problems. We also find that models fine-tuned on validated data still manage to learn a few functions removed during validation. Further, we release the synthetic dataset to seed future research in this area.
@@ 2 Related work
Tan et al. (2024) discusses the applications of LLMs in data annotation for classification tasks. Goel et al. (2023) demonstrates the use of LLMs in the medical domain, where they assist in labeling data with expert verification. Wang et al. (2024), Kim et al. (2024), and tang2024pdfchatannotator explore human-LLM collaborative approaches for annotation and verification. There has been no comparison of NL validation techniques on synthetic NL for NL2F.
@@ 3 Validating synthetic data
This validator asks the LLM to directly predict the output values $F(T)$ from $(\hat{U},T)$ and uses an element-wise row comparison to evaluate correctness. This approach leverages natural language to emulate the computation directly. It is inspired from the alternate task of output prediction discussed in Khatry et al. (2023)
@@
More details about the validators are provided in Appendix 8.3.
@@ 4 Experimental setup
We use two open (phi-2 (2B) and mistral-7b-instruct (7B)) and two closed-weight (gpt-35-turbo and gpt-4) models. phi-2 (8 $\times$ V100) and mistral (1 $\times$ A100) were fine-tuned for 10 and 15 epochs respectively. We selected the best checkpoint using validation loss. gpt-35 ($16$$\times$ A100) and gpt-4 ($24$$\times$ A100) were fine-tuned using the Azure API. mistral, gpt-35, gpt-4 were fine-tuned using LoRA Hu et al. (2021).
@@
We use the pass@$k$ metric Chen et al. (2021a) based on execution match of formula, were $k$ represents the number of predictions considered out of the total number of predictions provided. In our evaluation system, we generate $n=10$ predictions at temperature $0.6$ and compute pass@5 metric.
@@ 5 Results and Discussion
  • RQ4
    Can models finetuned on validated data learn the functions removed during validation?
@@ 5.2 RQ2: Effect on fine-tuning performance
Table 2 shows base model (few-shot) and fine-tuning performance on different subsets of data. For the smaller models, phi-2 and mistral, the performance increase with fine-tuning is more significant. With all models, a smaller, validated dataset yields better performance than raw data. $V_{P}$ yields the best performance on average with nearly half the size of raw data. gpt-4 improves only when fine-tuned on validated data. Surprisingly, gpt-35 without fine-tuning outperforms the fine-tuned version, likely due to differences in data distribution between training and testing benchmarks. Besides performance, fine-tuning with validated data also reduces training time significantly, as shown in Table 6. We see the performance on dataset created by the intersection of all validators (marked by $\cap$) is limited by the worst performing validator in each case.
Table 2: Performance comparison of the different models on SofSet Benchmark using pass@5 metric. Three out of the four models give best performance when fine-tuned on data validated by $V_{P}$.
FT on # ex phi-2 mistral gpt-35 gpt-4
Base 0 0.02 0.04 0.30 0.36
Raw 7833 0.08 0.16 0.29 0.32
$V_{O}$ 2266 0.06 0.17 0.28 0.35
$V_{P}$ 4095 0.10 0.14 0.30 0.41
$V_{C}$ 5246 0.09 0.14 0.30 0.40
$\cap$ 1607 0.06 0.13 0.27 0.34
@@
Table 3 compares the performance of fine-tuning on the accepted (& subsampled) and rejected ($\neg$) examples for each validator. We sub-sample the accepted sets to 2266—the number of examples in the smallest set ($V_{O}$). We observe that, despite the smaller size of the validated data subset (subsampled), it outperforms its larger invalidated (rejected) counterpart in most (11/12) comparisons. The only case where this not happens is for $V_{O}$ on gpt-4, likely due to the many functions (51) that were eliminated from the training data.
Table 3: Pairwise comparison of performance of sub-sampled ($\subset$) data from validated ($V$) against rejected ($\neg V$) examples. Results of pairs $(\subset V,\neg V)$ are marked in green if $(\subset V>\neg V)$, blue if $(\subset V=\neg V)$.
FT on # ex phi-2 mistral gpt-35 gpt-4
$V_{O}$ 2266 0.06 0.17 0.28 0.35
$\neg V_{O}$ 5567 0.05 0.16 0.27 0.35
$\subset{V_{C}}$ 2266 0.07 0.14 0.27 0.36
$\neg V_{C}$ 2587 0.04 0.12 0.26 0.34
$\subset V_{P}$ 2266 0.08 0.12 0.31 0.37
$\neg V_{P}$ 3738 0.05 0.11 0.24 0.32
@@ 5.3 RQ3: Analysing solved cases
Figure 3 shows properties of the solved cases (where atleast one prediction was correct) after fine-tuning different models on raw data and validated subsets. We see that fine-tuning on datasets with fewer unique functions still enables all models (except for mistral) to solve cases with more unique functions. The average function call count increases for validated subsets compared to the raw data, indicating more complex formulas are solved by models fine-tuned on validated data. For gpt-4 and gpt-35, average operator count also increases with fine-tuning on validated data.
@@
5.4 RQ4: Recovery of functions removed during validation
By analyzing the output generations of the fine-tuned models, we identify new functions that were not present in their base model predictions. Our results show that the models successfully learned some of the functions that had been removed during validation (see Table 4). This suggests that fine-tuning on a high-quality dataset allows the model to remember knowledge that it had learned during pre-training, without teaching it to hallucinate on potential mistakes in the synthetic data.
Table 4: Number of functions learned by different fine-tuned models that were removed during validation. Some examples include ’SUMIF’, ’TIME’, ’QUOTIENT’,’ROWS’,’AGGREGATE’
phi-2 mistral gpt-35 gpt-4
$V_{0}$ 1 2 2 3
$V_{P}$ 3 2 3 3
$V_{C}$ 1 1 2 2
@@ 6 Conclusion
We empirically evaluate the effect of automated validation of synthetic data using LLMs on the fine-tuning performance of derived-column NL-to-formula. We validate synthetic NL annotations with three surrogate tasks (classification, code generation in Python, and output prediction) and fine-tune different models on the examples accepted by each of these methods. In general, fine-tuning on smaller, validated datasets improves performance. Despite validation resulting in datasets with simpler formulas, that does not cause the fine-tuned models to only solve simpler problems. Models fine-tuned on validated data are able to recover some functions that were removed during validation.